# Many-Body Localization in the Age of Classical Computing

### v2

@@ X.2 Quasiperiodic potential
The discussion above concentrated on the random disorder, i.e., the situation in which the on-site potential strengths $h_{i}$ are independent and identically distributed random variables. Introduction of the so-called quasiperiodic potential is another way of breaking the translational invariance in many-body systems that leads to a similar phenomena of ergodicity breaking. One example of a quasiperiodic potential is
$h_{i}=W\cos(2\pi\eta j+\varphi),$ (91)
@@
An early study [294] considered the XXZ model with diagonal quasiperiodic potential given by (91), finding a crossover between the ETH and MBL regimes with the increase of the potential strength $W$. The initial claim was that the crossover gives rise to the MBL transition, which is smoothly connected, at weak interaction strength $\Delta$, to the Aubry-André localization threshold. This claim was disputed by [737], which demonstrated that transport properties discontinuously change from localization to diffusion upon switching on the interactions. Nevertheless, the strong spatial correlations between on-site potentials $h_{i}$ present in the quasiperiodic case eliminate the possibility of emergence of the large Griffiths rare regions of anomalously weak (or strong) disorder. This constitutes a major difference with respect to the MBL problem in the random disorder case. Several aspects of the dynamics in quasiperiodic potentials (and their generalizations) has been addressed over the years in diverse systems [439, 583, 316, 365, 55, 701, 173, 702, 402, 608, 10, 735, 689, 711, 667, 695, 498, 654]. It is worth observing that the potential (91) depends on two parameters $W$ and $\eta$. It turns out that the position of the ETH to MBL crossover, $W^{*}(L)$, strongly depends on the value of the incommensurability factor $\eta$ as numerically observed [173]. This has been partially explained in the single-particle case [199]. Namely, the Aubry-André transition (i.e., the transition in the absence of interactions, at $\Delta=0$) is determined by the distribution of the neighboring differences $\delta_{i}=h_{i+1}-h_{i}$ whose amplitude depends on ${\eta}$. Nevertheless, a complete understanding of the impact of the value of ${\eta}$, or other parameters specifying the quasiperiodic potential, c.f. [402, 9], on the ETH-MBL crossover is still lacking.
@@
The absence of Griffiths regions in the quasiperiodic systems translate into the features of the crossover between the ETH and MBL regimes. Recall that the cost function analysis of the mean gap ratio for the uniform random disorder, put forward in [691], yields a linear growth of the characteristic crossover disorder value, $W_{cr}\sim L$, suggesting a BKT type of the transition. A similar analysis, performed for the quasiperiodic case [30], also pointed towards the BKT transition, however, it also suggested $W_{cr}\sim\ln L$ rather than $W_{cr}\sim L$, thereby exhibiting a weaker dependence on the system size $L$. Extrapolation of the drift leads to $W_{cr}\approx 4$ for $L=50$. Interestingly, around this value of $W$, the density autocorrelation function $C(t)$, defined in Eq. (47), for the initial Néel state reveals slow oscillations that are in phase for a wide interval of system system sizes $L\in[12,200]$. As argued in [604] this may indicate a truly localized character of the dynamics. This suggests that a clear evidence for the MBL phase may be found in quasiperiodic systems at smaller time and length scales than in the case of random disorder. This expectation is further supported by the faster than exponential decay of the diffusion constant in the XXZ spin chain with quasiperiodic potential [502]. Finally, we note that the analysis of the MBL regime from the Fock-space perspective allows to distinguish quasiperiodic and random systems by their sample-to-sample fluctuations [229], similarly to the entropy measures [316] and level statistics [602].
@@ X.3 Other sources/types of disorder
Several other interesting types of disorder were used over the years to study aspects related to the MBL. One example, important from the experimental perspective, is realized by the speckle disorder used in the non-interacting case for demonstrating Anderson localization [75]. It has been also applied in an early attempt [334] to observe MBL in 3D experiment. One may also use a speckle disorder on top of the optical lattice to create a random diagonal potential with variable correlation length [405] which affects the properties of the ETH-MBL crossover [547, 586]. While in the latter case, the numerically observed properties may be traced back to the spatial correlations of the potential, [734] considered a problem of finding a distribution of the on-site potential that would enhance the MBL signatures in a maximal way.
@@
Another interesting possibility is to introduce disorder by selecting the values of the on-site fields $h_{i}$ from a discrete set. To that end, [222] proposed to use heavy, practically immobile atoms, interacting with light particles of a different species. In that case, the heavy particles, on the time scales on which their motion can be neglected, produce an effective random binary potential for the light particles. If the light particles are confined to 1D and are themselves not interacting - the model leads to Anderson localization. The interactions between light particles together with the effective binary potential may give rise to onset of an MBL regime. Nevertheless, even a slow motion of the heavy particles was argued to lead to thermalization of such a system at suffiently long time scales [464].
@@
Nonetheless, one may consider a model with binary quenched disorder for which the on-site potentials are chosen from a two element set, $h_{i}\in\{-W,W\}$. The model is attractive from the theoretical perspective, as the two possible values of disorder may be modelled by extending the Hilbert space, adding an auxiliary quantum 1/2-spin for each lattice site. Quantum evolution of the enlarged system, following the proposition of [467], reproduces, in a single run, the average over all $2^{L}$ realizations of the binary disorder, see also [613]. This enables calculation of the autocorrelation function $C(t)$, for the initial Néel state, directly in the thermodynamic limit $L\to\infty$ [639, 194]. However, the time evolution in this approach is followed up to a relatively short time $t=30$. Hence, due to the observed oscillations and the decay of $C(t)$ the extrapolation of the results to the limit $t\to\infty$, relevant for the MBL phase, is not feasible. Exact diagonalization results of the XXZ spin chain with binary disorder leads to similar values of the exponent $\nu$ as for the typically studied unform random disorder [194], suggesting similarities between the finite size drifts in the two types of the model. Let us note that other discrete disorders may be studied in the context of the MBL transition [298]. For uncorrelated discrete disorders with more than 2 values, the ETH-MBL crossover in the XXZ spin chain shows a large degree of universality, at least for the system sizes accessible to exact diagonalization studies. [298] shows that various ergodicity breaking indicators $\overline{X}$, e.g., the gap ratio $\overline{r}$, plotted as a function of the variance of the on-site potential $h_{i}$ collapse for different types of disorder.
@@
All the models considered so far, with the exception of the PXP model, have one common feature. In the absence of interactions the models reduce to disordered single-particle problems in 1D, which are Anderson localized. In fact, the proximity of the Anderson localized limit provides the usual avenue for studies of the dynamics of strongly disordered many-body systems, as we argued in Sec. VII.3.1. Nevertheless, can we observe signatures of MBL when the system is delocalized in the absence of interactions? Such a problem was studied for bosons in a lattice with random on-site interactions [591, 592, 601]. In that case, $\mu_{i}=0$ and the on-site interaction term in (89) is modified as $\frac{U}{2}\sum_{i=1}^{L}\hat{n}_{i}(\hat{n}_{i}-1)\to\frac{1}{2}\sum_{i=1}^{L}U_{i}\hat{n}_{i}(\hat{n}_{i}-1)$, where $U_{i}$ is a site-dependent interaction amplitude, drawn independently at each site from a specified distribution. Independently, similar problems were considered for spinful fermions with random on-site interactions [45] as well as for spinless fermions [371]. In the absence of interactions, $U_{i}=0$, the systems are delocalized with Bloch waves forming the eigenstates. Turning on random interactions, $U_{i}>0$, which for ultracold bosons can be realized by placing the system close to a wave-guide [231], may lead to the MBL that clearly has no localized single-particle limit. Despite that, the characteristics of MBL regime observed both in quantities derived from eigenvalues and eigenvectors (e.g., the gap ratio), as well as features of time dynamics, e.g., the density autocorrelation function $C(t)$, show significant similarity with the standard disordered spin chains. Another model with similar properties was considered in [402]. Here, particles were considered on a Fibonacci chain and the non-interacting case revealed multifractal eigenstates. Better understanding of finite-size drifts in such models, as well as identification of their dynamical features distinct from the standard spin chains, remain interesting avenues for further research.
@@
Apart for the on-site disorder, one may also encounter disorder in tunnelling terms in the models in Eqs. (7), (89) and (88). Ground state physics of random-bond Heisenberg spin, for which $J_{i}$ in (7) are taken to be random variables, was studied with the renormalization group [400, 154, 205]. These approaches were extended to the non-equilibrium systems [687, 688] and the considerations were followed by [668, 539]. Notably, the random bond XXZ spin chain at $\Delta=1$ has an SU(2) symmetry, which was argued to be incompatible with the MBL phase [494, 512]. In spite of this, the random bond Heisenberg spin chain hosts a wide non-ergodic regime [513], and a similar wide non-ergodic regime is found in the SU(3) symmetric spin chains [147]. Better understanding of the finite size drifts at this crossover could allow us to form more precise formulations about the system in which the MBL phase may be expected to occur. A similar class of models with bond disorder has been recently studied theoretically in the context of chains of Rydberg atoms [605, 566, 638].
@@ X.4 Floquet MBL
The advantage of studying non-equlibrium many-body problems in Floquet systems was recognized by Tomaž Prosen long before the MBL era [505, 506, 507]. Remarkably, the Floquet systems became also the first system allowing for an analytical demonstration of quantum chaos in many-body systems [66, 336, 108, 65] [^39] Soon, it was realized that so-called dual-unitarity [67] was the property allowing for analytical insights..
@@
$U_{F}=e^{-ig\sum_{j=1}^{L}\sigma^{x}_{j}}e^{-i\sum_{j=1}^{L}(J\sigma^{z}_{j}\sigma^{z}_{j+1}+h_{j}\sigma^{z}_{j})},$ (92)
where $\sigma^{x,y,z}_{j}$ are Pauli operators, $h_{j}\in[0,2\pi]$ are independent and uniformly distributed random variables, and periodic boundary conditions are assumed. Setting $g=J=1/W$, the parameter $W$ plays the role of the disorder strength in the system. Solving the eigenproblem $U_{F}\ket{\psi_{n}}=e^{i\phi_{n}}\ket{\psi_{n}}$ yields the eigenphases $\phi_{n}\in[0,2\pi]$ and eigenvectors. The eigenproblem can be solved with full exact diagonalization or with an approach relying on polynomial filter approach [385] analogous to the POLFED algorithm, as described in Sec. V.1. The polynomial filter approach provides access to system sizes up to $L=20$ on present-day supercomputers, significantly beyond the scope of full exact diagonalization algorithms. The access to larger system sizes provides means to uncover quantitative changes in finite size drifts at the ETH-MBL crossover in the Kicked Ising model.
In the ergodic regime, the spectral properties of $U_{F}$ are modelled by Circular Orthogonal Ensemble of random matrices [258]. Hence, the ETH-MBL crossover in the Kicked Ising model can be investigated similarly to the autonomous [^40] In this review an autonomous system is a many-body quantum system with time-independent Hamiltonian. disordered spin chains, see Sec. VI.1. Consequently, [596] analyze the crossover using the system size dependent disorder strengths $W^{T}(L)$ and $W^{*}(L)$. The boundary of the ergodic regime $W^{T}(L)$ scales linearly with system size, $W^{T}(L)\propto L$, for $L\leq 15$, as in the autonomous disordered XXZ spin chains. Notably, this trend is replaced by a sub-linear growth at $L\geq 15$, consistent with the possibility of a transition to the MBL phase at a sufficiently strong disorder. The crossing point $W^{*}(L)$ is drifting with $L$, however, the drift slows down with the increase of system size, and $W^{*}(L)$ is well fitted by a first order polynomial in $1/L$. Moreover, extrapolation of the linear scaling of $W^{T}(L)$, found in small systems, crosses the extrapolation of $W^{*}(L)$ at $L\approx 28$. The latter length scale is significantly smaller than $L\approx 50$ found for the disordered XXZ spin chain. The finite size drifts observed for several variants of kicked spin chains [596] are similar to those in the Kicked Ising model, and analogous to the finite size drifts found in the Anderson model on random regular graphs [595, 671].
@@ X.5 Long-range interactions
A systematic analysis of non-equilibrium physics in models with power-law decaying couplings was carried out by [715]. The Hamiltonian considered reads:
@@
Since then, a number of works considered different aspects of MBL-like behaviour in disordered long-range interacting systems, e.g., [267, 369, 255, 610, 441, 656, 543, 161, 436, 533, 84, 404, 555, 329, 167, 719, 652]. Several signatures of MBL regime were discussed, and the decay of LIOMs in systems with long-range interactions was proposed to be algebraic as opposed to the exponential decay in short-range MBL systems. One of the characteristic features of the MBL regime in the presence of long-range (algebraic) interactions turns out to be algebraic (and not logarithmic, as expected for short-range interactions) growth of the entanglement entropy [543, 161]. Such an algebraic growth persists from deep inside the MBL regime down to the crossover to the ergodic regime. In fact, it has been numerically found [167] that the growth $S\sim t^{\gamma}$, with $\gamma=1/3$, corresponds to the characteristic disorder value designating the crossover in a broad range of the values of exponents $a$ and $b$ describing the interaction range in (93). One may also consider a model of infinite-range, i.e., all-to-all, interactions as those mediated by photons in the cavity [589, 347, 113]. In that case, the entanglement entropy growth in the MBL regime is also well fitted by an algebraic dependence on time [589]. Quantum dots provide another class of models with all-to-all interactions interesting from the perspective of ergodicity breaking phenomena, see e.g. [272] and references therein.
The described results regarding the non-equilibrium physics of systems with long-range interactions were obtained numerically for systems of size up to $L=20$, or for larger systems with tensor networks but over short evolution times. If the interactions are decaying according to a power-law or slower, the avalanche scenario, see Sec. IX, predicts the absence of MBL phase. Indeed, if the power-law decay of matrix elements was included in Eq. (71), the balance between the decay of matrix element and the growth of the ergodic bath’s density of states could not be attained, resulting in the thermalization of the system. Nevertheless, long-range interactions, contrary to naive expectations, may lead to slow growths of entanglement [466, 362], illustrating the need of a better understanding of microscopic origins of ergodicity in the presence of long-range couplings. Solvable all-to-all Sachdev-Ye-Kitaev (SYK) models [130] provide a reference point for these questions. Further quantitative explorations of the drifts at the ETH-MBL crossover in the presence of long-range interactions are needed to better understand its relation to the crossover found in the disordered short-range systems.

### v3

@@ X.2 Quasiperiodic potential
The discussion above concentrated on the random disorder, i.e., on the situation in which the on-site potential strengths $h_{i}$ are independent and identically distributed random variables. Introduction of the so-called quasiperiodic potential is another way of breaking the translational invariance in many-body systems that leads to a similar phenomena of ergodicity breaking. One example of a quasiperiodic potential is
$h_{j}=W\cos(2\pi\eta j+\varphi),$ (91)
@@
An early study [319] considered the XXZ model with diagonal quasiperiodic potential given by (91), finding a crossover between the ETH and MBL regimes with the increase of the potential strength $W$. The initial claim was that the crossover gives rise to the MBL transition, which is smoothly connected, at weak interaction strength $\Delta$, to the Aubry-André localization threshold. This claim was disputed by [794], which demonstrated that transport properties discontinuously change from localization to diffusion upon switching on the interactions. Nevertheless, the strong spatial correlations between on-site potentials $h_{i}$, present in the quasiperiodic case, eliminate the possibility of emergence of the large Griffiths regions of anomalously weak (or strong) disorder. This constitutes a major difference with respect to the MBL problem in the random disorder case. Several aspects of the dynamics in quasiperiodic potentials (and their generalizations) has been addressed over the years in diverse systems [477, 630, 342, 397, 58, 757, 182, 758, 438, 657, 10, 792, 745, 767, 720, 751, 539, 706]. It is worth observing that the potential (91) depends on two parameters $W$ and $\eta$. It turns out that the position of the ETH to MBL crossover, $W^{*}(L)$, strongly depends on the value of the incommensurability factor $\eta$ as numerically observed [182]. This has been partially explained in the single-particle case [209]. Namely, the Aubry-André transition (i.e., the transition in the absence of interactions, at $\Delta=0$) is determined by the distribution of the neighboring differences $\delta_{i}=h_{i+1}-h_{i}$ whose amplitude depends on ${\eta}$. Nevertheless, a complete understanding of the impact of the value of ${\eta}$, or other parameters specifying the quasiperiodic potential, c.f. [438, 9], on the ETH-MBL crossover is still lacking.
@@
The absence of Griffiths regions in the quasiperiodic systems translate into the features of the crossover between the ETH and MBL regimes. Recall that the cost function analysis of the mean gap ratio for the uniform random disorder, put forward in [747], yields a linear growth of the characteristic crossover disorder value, $W_{cr}\sim L$, suggesting a BKT type of the transition. A similar analysis, performed for the quasiperiodic case [32], also pointed towards the BKT transition, however, it also suggested $W_{cr}\sim\ln L$ rather than $W_{cr}\sim L$, thereby exhibiting a weaker dependence on the system size $L$. Extrapolation of the drift leads to $W_{cr}\approx 4$ for $L=50$. Interestingly, around this value of $W$, the density autocorrelation function $C(t)$, defined in Eq. (47), for the initial Néel state reveals slow oscillations that are in phase for a wide interval of system system sizes $L\in[12,200]$. As argued in [653], this may indicate a truly localized character of the dynamics. This suggests that a clear evidence for the MBL phase may be found in quasiperiodic systems at smaller time and length scales than in the case of random disorder. This expectation is further supported by the faster than exponential decay of the diffusion constant in the XXZ spin chain with quasiperiodic potential [543]. Finally, we note that the analysis of the MBL regime from the Fock-space perspective allows to distinguish quasiperiodic and random systems by their sample-to-sample fluctuations [245], similarly to the entropy measures [342] and level statistics [651].
@@ X.3 Other sources/types of disorder
Several other interesting types of disorder were used over the years to study aspects related to the MBL. One example, important from the experimental perspective, is realized by the speckle disorder used in the non-interacting case for demonstrating Anderson localization [78]. It has been also applied in an early attempt [362] to observe MBL in a 3D experiment. One may also use a speckle disorder on top of the optical lattice to create a random diagonal potential with variable correlation length [441], which affects the properties of the ETH-MBL crossover [593, 633]. While in the latter case, the numerically observed properties may be traced back to the spatial correlations of the potential, [791] considered a problem of finding a distribution of the on-site potential that would enhance the MBL signatures in a maximal way.
@@
Another interesting possibility is to introduce disorder by selecting the values of the on-site fields $h_{i}$ from a discrete set. To that end, [237] proposed to use heavy, practically immobile atoms, interacting with light particles of a different species. In that case, the heavy particles, on the time scales on which their motion can be neglected, produce an effective random binary potential for the light particles. If the light particles are confined to one dimension and are themselves not interacting, the model leads to Anderson localization. The interactions between light particles together with the effective binary potential may give rise to the onset of an MBL regime. Nevertheless, even a slow motion of the heavy particles was argued to lead to thermalization of such a system at suffiently long time scales [502].
@@
Nonetheless, one may consider a model with binary quenched disorder for which the on-site potentials are chosen from a two element set, $h_{i}\in\{-W,W\}$. The model is attractive from the theoretical perspective, as the two possible values of disorder may be modelled by extending the Hilbert space, adding an auxiliary quantum 1/2-spin for each lattice site. Quantum evolution of the enlarged system, following the proposition of [505], reproduces, in a single run, the average over all $2^{L}$ realizations of the binary disorder, see also [663]. This enables calculation of the autocorrelation function $C(t)$, for the initial Néel state, directly in the thermodynamic limit $L\to\infty$ [690, 204]. However, the time evolution in this approach is followed up to a relatively short time $t=30$. Hence, due to the observed oscillations and the decay of $C(t)$, the extrapolation of the results to the limit $t\to\infty$, relevant for the MBL phase, is not feasible. Exact diagonalization results of the XXZ spin chain with binary disorder leads to similar values of the exponent $\nu$ as for the typically studied unform random disorder [204], suggesting similarities between the finite size drifts in the two types of the model. Let us note that other discrete disorders may be studied in the context of the MBL transition [323]. For uncorrelated discrete disorders with more than 2 values, the ETH-MBL crossover in the XXZ spin chain shows a large degree of universality, at least for the system sizes accessible to exact diagonalization studies. [323] showed that various ergodicity breaking indicators $\overline{X}$, e.g., the gap ratio $\overline{r}$, plotted as a function of the variance of the on-site potential $h_{i}$, collapse for different types of disorder.
@@
All the models considered so far, with the exception of the PXP model, have one common feature. In the absence of interactions the models reduce to disordered single-particle problems in one dimension, which are Anderson localized. In fact, the proximity of the Anderson localized limit provides the usual avenue for studies of the dynamics of strongly disordered many-body systems, as we argued in Sec. VII.3.1. Nevertheless, can we observe signatures of MBL when the system is delocalized in the absence of interactions? Such a problem was studied for bosons in a lattice with random on-site interactions [640, 641, 650]. In that case, $\mu_{i}=0$ and the on-site interaction term in (89) is modified as $\frac{U}{2}\sum_{i=1}^{L}\hat{n}_{i}(\hat{n}_{i}-1)\to\frac{1}{2}\sum_{i=1}^{L}U_{i}\hat{n}_{i}(\hat{n}_{i}-1)$, where $U_{i}$ is a site-dependent interaction amplitude, drawn independently at each site from a specified distribution. Independently, similar problems were considered for spinful fermions with random on-site interactions [48] as well as for spinless fermions [405]. In the absence of interactions, $U_{i}=0$, the systems are delocalized with Bloch waves forming the eigenstates. Turning on random interactions, $U_{i}>0$, which for ultracold bosons can be realized by placing the system close to a wave-guide [247], may lead to the MBL that clearly has no localized single-particle limit. Despite that, the characteristics of MBL regime observed both in quantities derived from eigenvalues and eigenvectors (e.g., the gap ratio), as well as features of time dynamics, e.g., the density autocorrelation function $C(t)$, show significant similarity with the standard disordered spin chains. Another model with similar properties was considered in [438]. Here, particles were considered on a Fibonacci chain and the non-interacting case revealed multifractal eigenstates. Better understanding of finite-size drifts in such models, as well as identification of their dynamical features distinct from the standard spin chains, remain interesting avenues for further research.
@@
Apart for the on-site disorder, one may also encounter disorder in tunnelling terms in the models in Eqs. (7), (89) and (88). The standard formulation of Lieb-Robinson bounds [407] is insufficient to capture the complexity of the dynamics induced by "weak links", i.e., bonds for which $J_{i}$ in (7) is anomalously small. Instead, capturing the rich dynamics of such models requires a distinction between the bounds that hold for all sites of the chain and bounds that hold for a subsequence of sites, as shown in [45]. Ground state physics of random-bond Heisenberg spin, for which $J_{i}$ are taken to be random variables, was studied using the renormalization group [436, 160, 218]. These approaches were extended to the non-equilibrium systems [742, 743] and were followed by [721, 585]. Notably, the random bond XXZ spin chain at $\Delta=1$ has an SU(2) symmetry, which was argued to be incompatible with the MBL phase [534, 553]. In spite of this, the random bond Heisenberg spin chain hosts a wide non-ergodic regime [369, 554], and a similar wide non-ergodic regime is found in the SU(3) symmetric spin chains [152]. Better understanding of the finite size drifts at this crossover could allow us to form more precise formulations about the system in which the MBL phase may be expected to occur. A similar class of models with bond disorder has been recently studied theoretically in the context of chains of Rydberg atoms [654, 613, 689].
@@ X.4 Floquet MBL
The advantage of studying non-equlibrium many-body problems in Floquet systems was recognized by Tomaž Prosen long before the MBL era [546, 547, 548]. Remarkably, the Floquet systems became also the first system allowing for an analytical demonstration of quantum chaos in many-body systems [69, 364, 112, 68] [^49] Soon, it was realized [70] that so-called dual-unitarity, introduced in [14], was the property allowing for analytical insights..
@@
$\hat{U}_{F}=e^{-ig\sum_{j=1}^{L}\hat{\sigma}^{x}_{j}}e^{-i\sum_{j=1}^{L}(J\hat{\sigma}^{z}_{j}\hat{\sigma}^{z}_{j+1}+h_{j}\hat{\sigma}^{z}_{j})},$ (92)
where $\hat{\sigma}^{x,y,z}_{j}$ are Pauli operators, $h_{j}\in[0,2\pi]$ are independent and uniformly distributed random variables, and periodic boundary conditions are assumed. Setting $g=J=1/W$, the parameter $W$ plays the role of the disorder strength in the system. Solving the eigenproblem $\hat{U}_{F}\ket{\psi_{n}}=e^{i\phi_{n}}\ket{\psi_{n}}$ yields the eigenphases $\phi_{n}\in[0,2\pi]$ and eigenvectors. The eigenproblem can be solved with full exact diagonalization or with an approach relying on polynomial filter approach [421] analogous to the POLFED algorithm, as described in Sec. V.1. The polynomial filter approach provides access to system sizes up to $L=20$ on present-day supercomputers, significantly beyond the scope of full exact diagonalization algorithms. The access to larger system sizes provides means to uncover quantitative changes in finite size drifts at the ETH-MBL crossover in the Kicked Ising model.
In the ergodic regime of isolated periodically driven interacting systems [154], the spectral properties are modelled by Circular Orthogonal Ensemble of random matrices [274]. Hence, the ETH-MBL crossover in the Kicked Ising model can be investigated similarly to the autonomous [^50] In this review an autonomous system is a many-body quantum system with time-independent Hamiltonian. disordered spin chains, see Sec. VI.1. Consequently, [645] analyzed the crossover using the system size dependent disorder strengths $W^{T}(L)$ and $W^{*}(L)$. The boundary of the ergodic regime $W^{T}(L)$ scales linearly with system size, $W^{T}(L)\propto L$, for $L\leq 15$, as in the autonomous disordered XXZ spin chains. Notably, this trend is replaced by a sub-linear growth at $L\geq 15$, consistent with the possibility of a transition to the MBL phase at a sufficiently strong disorder. The crossing point $W^{*}(L)$ is drifting with $L$, however, the drifts slows down with the increase of system size, and $W^{*}(L)$ is well fitted by a first order polynomial in $1/L$. Moreover, extrapolation of the linear scaling of $W^{T}(L)$, found in small systems, crosses the extrapolation of $W^{*}(L)$ at $L\approx 28$. The latter length scale is significantly smaller than $L\approx 50$ found for the disordered XXZ spin chain. The finite size drifts observed for several variants of kicked spin chains [645] are similar to those in the Kicked Ising model, and also similar to the finite size drifts found in the Anderson model on random regular graphs [644, 725].
@@ X.5 Long-range interactions
A systematic analysis of non-equilibrium physics in models with power-law decaying couplings was carried our by [771]. The Hamiltonian considered reads:
@@
Since then, a number of works considered different aspects of MBL-like behaviour in disordered long-range interacting systems, e.g., [286, 402, 271, 660, 479, 708, 589, 168, 474, 579, 87, 440, 602, 357, 175, 775, 704]. Several signatures of MBL regime were discussed, and the decay of LIOMs in systems with long-range interactions was proposed to be algebraic as opposed to the exponential decay in short-range MBL systems. One of the characteristic features of the MBL regime in the presence of long-range (algebraic) interactions turns out to be algebraic (and not logarithmic, as expected for short-range interactions) growth of the entanglement entropy [589, 168]. Such an algebraic growth persists from deep inside the MBL regime down to the crossover to the ergodic regime. In fact, it has been numerically found [175] that the growth $S\sim t^{\gamma}$, with $\gamma=1/3$, corresponds to the characteristic disorder value designating the crossover in a broad range of the values of exponents $a$ and $b$ describing the interaction range in (93). One may also consider a model of infinite-range, i.e., all-to-all, interactions as those mediated by photons in the cavity [638, 376, 117, 238]. In that case, the entanglement entropy growth in the MBL regime is also well fitted by an algebraic dependence on time [638]. Quantum dots provide another class of models with all-to-all interactions interesting from the perspective of ergodicity breaking phenomena, see e.g. [291] and references therein.
The described results regarding the non-equilibrium physics of systems with long-range interactions were obtained numerically for systems of size up to $L=20$, or for larger systems with tensor networks but over short evolution times. If the interactions are decaying according to a power-law or slower, the avalanche scenario, see Sec. IX, predicts the absence of MBL phase. Indeed, if the power-law decay of matrix elements was included in Eq. (71), the balance between the decay of matrix element and the growth of the ergodic bath’s density of states could not be attained, resulting in the thermalization of the system. Nevertheless, long-range interactions, contrary to naive expectations, may lead to slow growths of entanglement [504, 394], illustrating the need of a better understanding of microscopic origins of ergodicity in the presence of long-range couplings [172]. Solvable all-to-all Sachdev-Ye-Kitaev (SYK) models [135] provide a reference point for these questions. Further quantitative explorations of the drifts at the ETH-MBL crossover in the presence of long-range interactions are needed to better understand its relation to the crossover found in the disordered short-range systems.